Investors in Ranke Electric’s stock require a return of 7.3%. If the company simply earns the cost of capital on its new investments, what is the stock’s P/E? (Round your answer to 2 decimal place.)

Stock P/E Ratio
Answer #1

Return = 1/(P/E ratio)

P/E ratio=1/return
=1/0.073
=13.70

Hence P/E for the stock is 13.70